# Simplified and Generalized Masked Diffusion for Discrete Data

### v1

@@ Abstract
Masked (or absorbing) diffusion is actively explored as an alternative to autoregressive models for generative modeling of discrete data. However, existing work in this area has been hindered by unnecessarily complex model formulations and unclear relationships between different perspectives, leading to suboptimal parameterization, training objectives, and ad hoc adjustments to counteract these issues. In this work, we aim to provide a simple and general framework that unlocks the full potential of masked diffusion models. We show that the continuous-time variational objective of masked diffusion models is a simple weighted integral of cross-entropy losses. Our framework also enables training generalized masked diffusion models with state-dependent masking schedules. When evaluated by perplexity, our models trained on OpenWebText surpass prior diffusion language models at GPT-2 scale and demonstrate superior performance on 4 out of 5 zero-shot language modeling tasks. Furthermore, our models vastly outperform previous discrete diffusion models on pixel-level image modeling, achieving 2.78 (CIFAR-10) and 3.42 (ImageNet 64$\times$64) bits per dimension that are comparable or better than autoregressive models of similar sizes.
@@ 1 Introduction
  • By using elementary arguments, we establish several properties for the forward process induced by this model and its corresponding time reversal, improving our understanding of this model class.
@@
  • Finally, based on our simplified masked diffusion formulation, we propose a generalized masked diffusion model that allows state-dependent masking schedules. This generalized masked diffusion model further improves predictive performance on test likelihoods.
@@ 2 Masked Diffusion
The forward process is defined as a Markovian sequence of discrete random variables $x_{t}$ indexed by time $t$, where $t$ runs from 0 to 1. Throughout the work, we abuse the notation such that $x_{t}$ can be either an integer or its corresponding one-hot vector, whenever it is clear from the context. We divide $[0,1]$ into $T$ intervals, and let $s(i)=(i-1)/T$, $t(i)=i/T$. Following Austin et al. [14], the state transition between $[s(i),t(i)]$ is determined by a transition matrix of size $m\times m$: $Q_{i}=(1-\beta_{i})I+\beta_{i}\mathbf{1}e_{m}^{\top},$ where $\mathbf{1}$ is an all-one vector of size $m+1$, $e_{m}$ represents a one-hot vector with the $m$-th element being 1. Each entry $[Q_{i}]_{jk}$ denotes the probability of transition from the state $j$ to the state $k$:
@@
This means, with probability $1-\beta_{i}$, $x_{t(i)}=x_{s(i)}$, otherwise it jumps to the mask state. Given the above transition matrix, the marginal distribution at time $t(i)$ given $x_{0}$ is
@@
We can define a continuous-time forward process by taking a limit of the above discrete-time process. We first specify a continuous function $\beta(t)$ such that $\beta_{i}=\beta(t(i))/T$. We then let $T\to\infty$ in the discrete-time process and compute the limit of $\bar{Q}_{i}$ (proved in Austin et al. 14, Appendix A.6, see also Appendix A and Appendix B) as
@@
Figure 1: Masking schedules in the literature: (Left) $\alpha_{t}$; (Right) weight of the cross-entropy loss w.r.t. $t$; Equations for these schedules are given in Table 4 in Appendix.
@@
From the definition of $\alpha_{t}$, we have that $\alpha_{0}=1$. And similar to the discrete-time formulation, we would like $\alpha_{1}$ be zero or very close to zero. We provide a summary of masking schedules from literature that satisfy these properties in Figure 1. The linear schedule was proposed in Sohl-Dickstein et al. [1] for binary variables and then re-derived by Austin et al. [14] from mutual information for discrete-time models. The geometric schedule $\alpha_{t}$ is plotted for $\bar{\beta}_{\text{min}}=10^{-5}$ and $\bar{\beta}_{\text{max}}=20$. It was first used for continuous diffusions [3] and then for discrete by Lou et al. [32]. The cosine schedule was originally proposed in MaskGIT [30], an iterative unmasking generative model inspired by diffusion. This schedule has the property of slowing down the unmasking process at the beginning of the reverse generation. Aligning with their observation, we find that this results in a lower chance of conflicting tokens being unmasked simultaneously at the start of generation, thereby enhancing the overall generation quality.
@@
The analytic property of our forward process allows to compute many quantities of interest in closed form. One such quantity frequently used in diffusion models is the time reversal of the forward process given $x_{0}$: $q(x_{s}|x_{t},x_{0})$ for $s\leq t$. We derive it in Appendix C as
@@ 3 Model and Objective
For a discrete-time masked diffusion process, we define our generative model by approximately reversing the forward transitions using a backward model $p_{\theta}(x_{s}|x_{t})$. One way to define this model is
@@
This is known as mean-parameterization since it leverages a prediction model for the mean of $x_{0}$. Besides $p_{\theta}(x_{s}|x_{t})$, we also need to specify $p(x_{0}|x_{t(1)})$ and the prior distribution $p(x_{t(T)})=p(x_{1})$. Following the practice in continuous diffusion models [33], we choose $p(x_{0}|x_{t(1)})\propto q(x_{t(1)}|x_{0})$. And since $q(x_{1}|x_{0})\approx\delta_{x_{1},m}$ for any $x_{0}$ as $\alpha_{1}\approx 0$, we set $p(x_{1})\approx\delta_{x_{1},m}$, see Appendix E.
@@
where $\mathcal{L}_{T}=\sum_{i=2}^{T}\mathbb{E}_{q(x_{t(i)}|x_{0})}[\mathrm{KL}({q(x_{s(i)}|x_{t(i)},x_{0})}\|{p_{\theta}(x_{s(i)}|x_{t(i)})})]$. For the above choices of the prior distribution, the term $\mathrm{KL}({q(x_{1}|x_{0})}\|{p(x_{1})})$ becomes zero. Under the backward model (2), the KL divergence terms in $\mathcal{L}_{T}$ becomes (proof in Appendix D)
@@
which is a simple cross-entropy loss between the predicted logits and the clean data. In Appendix D, we show that $\mathcal{L}_{T}$ is a Riemann sum and is lower bounded by the corresponding continuous integral:
@@
For continuous state-space diffusions, it was shown that the ELBO is invariant to the noise schedule, except for the signal-to-noise ratio (SNR) at its endpoints [33]. We establish here a similar result for discrete diffusions. Consider choosing $\alpha_{t}=\sigma(\lambda_{t})$, where $\sigma$ represents the sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$. In this context, the log-SNR is defined by $\lambda_{t}=\log\frac{\alpha_{t}}{1-\alpha_{t}}=\textup{log-SNR}(t)$. By making a change of variables in (4) to make everything a function of the log-SNR, we obtain
$\mathcal{L}_{\infty}=\int_{\lambda_{t(1)}}^{\lambda_{1}}\sigma(\lambda)x_{0}^{\top}\mathbb{E}_{\tilde{q}(x_{\lambda}|x_{0})}\left[\delta_{x_{\lambda},m}\log\tilde{\mu}_{\theta}(x_{\lambda},\lambda)\right]\mathrm{d}\lambda.$
@@
In the previous sections, $x_{t}$ was assumed to be a single discrete token. To extend the method to multidimensional data, let $x_{t}$ be now a sequence $(x_{t}^{(1)},x_{t}^{(2)},\dots,x_{t}^{(N)})$, where each element $x_{t}^{(n)}$ represents a discrete token. We select a forward process which factorizes across all $N$ tokens: $q(x_{t}|x_{s})=\prod_{n=1}^{N}q(x_{t}^{(n)}|x_{s}^{(n)})$. As a result, the forward marginals $q(x_{t}|x_{0})$ and reversal $q(x_{s}|x_{t},x_{0})$ also factorize. In this case, we define the backward model as $p_{\theta}(x_{s}|x_{t})\triangleq\prod_{n=1}^{N}q(x_{s}^{(n)}|x_{t}^{(n)},\mu_{\theta}^{(n)}(x_{t},t))$, where $\mu_{\theta}(x_{t},t)$ is a neural network that takes the full $N$ tokens as input and outputs $N$ probability vectors. The $n$-th output $\mu_{\theta}^{(n)}(x_{t},t)$ is a prediction model for $\mathbb{E}[x_{0}^{(n)}|x_{t}]$, the mean value of the $n$-th token. Repeating above derivations gives
@@ 4 Relation to Existing Work
To show the connection with the CTMC view presented in Austin et al. [14], Campbell et al. [29], we can write out the forward and backward masked diffusion using CTMC machinery. To see this, for a short time $\Delta t$, given $x_{0}$, the Taylor expansions of our forward and reverse transition matrices at $t$ are
@@
$\bar{R}^{x_{0}}(t,t-\Delta t)$ $=I+R^{x_{0}}(t)\Delta t+o(\Delta t)\text{\quad for\quad}R^{x_{0}}(t)\triangleq-\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top}.$ (7)
where $Q(t)$ and $R^{x_{0}}(t)$ are known as the transition rate matrices. Austin et al. [14] derived the same $Q(t)$ in Appendix A.6 of their paper. However, they did not explore the reverse process or a continuous-time objective. Campbell et al. [29] obtained the continuous-time ELBO up to a constant from the CTMC view. They established the following alternative expression for the ELBO:
where C is a constant independent of $\theta$ and $q_{t|0}(k|x_{0}):=q(x_{t}=k|x_{0})$ for any $k$. We use the shorthand $R_{\theta}(t)_{kj}$ to denote the approximate backward transition rate from the state $k$ to $j$ obtained by substituting our prediction model $\mu_{\theta}(k,t)$ for $x_{0}$ in $R^{x_{0}}(t)_{kj}$. In Section F.1, we show how to recover this expression from by separating out a constant from our ELBO expression (4) and applying a discrete “integration-by-part”. Campbell et al. [29] used (8) as the training loss. A key limitation of this loss is that it requires $m$ evaluations of the prediction model $\mu_{\theta}(\cdot,t)$ to compute the inner summation term. To circumvent this computational burden, they proposed using a doubly stochastic estimate. However, this leads to significantly higher variance compared to the analytic cross-entropy (4) which only requires one pass of $\mu_{\theta}(\cdot,t)$. Please refer to Section F.2 for more details.
@@
While so far we used a prediction model $\mu_{\theta}(x_{t},t)$ for the mean of clean data given $x_{t}$ (i.e., mean parameterization), one can choose other ways of parameterizing the reverse model. Benton et al. [35], Lou et al. [32] proposed to parameterize the discrete “score” $s(x_{t},t)_{j}\triangleq\frac{q_{t}(j)}{q_{t}(x_{t})}$ and introduced the following score-based loss for discrete diffusions:
$\mathcal{L}_{s_{\theta}}\!=\!\int_{0}^{1}\!\mathbb{E}_{q_{t|0}(k|x_{0})}\Big{[}\sum_{j\neq k}Q(t)_{jk}\Big{(}s_{\theta}(k,t)_{j}\!-\!\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\log s_{\theta}(k,t)_{j}\!+\!\psi\Big{(}\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\Big{)}\Big{)}\Big{]}\mathrm{d}t,$ (9)
where $\psi(y)\triangleq y\log y-y$. In Section F.3, we provide an alternative derivation of (9) which is simpler. We show the link between score and mean parameterizations through the following proposition.
Proposition 1.
Let $q_{t}$ be the marginal distribution of the masked diffusion defined in Section 2 at time $t$. The discrete score $s(x_{t},t)_{j}=\frac{q_{t}(j)}{q_{t}(x_{t})}$ for a mask state $x_{t}=m$ and $j\neq m$ can be expressed as
@@
Proposition 1 (proved in Section F.3) implies that a reasonable score model for a mask state is
$s_{\theta}(m,t)_{j}=\frac{\alpha_{t}}{1-\alpha_{t}}\mu_{\theta}(m,t).$ (11)
Indeed, plugging (11) into the score-based loss (9) recovers our objective (4). In Lou et al. [32] the score is parameterized as a neural network without enforcing the constraint in (10). This means the learned backward model can be incompatible with the forward process. We find that our parameterization, which enforces the constraint, leads to more stable training and better results.
@@
Training of MaskGIT [30] follows the steps: (a) Sample $t\in[0,1]$. (b) Given a mask scheduling function $\gamma(t)$, sample $\gamma(t)N$ tokens to place masks. (c) For the data $x_{0}\in\mathbb{R}^{(m+1)\times N}$ and the masked token state $x_{t}\in\mathbb{R}^{(m+1)\times N}$, minimize the negative log-likelihood
From (10) we know that $\frac{\sum_{j\neq m}q_{t|0}(j|x_{0})}{q_{t|0}(m|x_{0})}=\frac{\alpha_{t}}{1-\alpha_{t}}$ thus $q_{t|0}(m|x_{0})=1-\alpha_{t}$. Therefore, when we set the mask scheduling function as $\gamma(t)=1-\alpha_{t}$ we can recover the loss in (5) without the $\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}$ weighting. One might be interested in whether the uniform weighting can be recovered by selecting an appropriate schedule $\alpha_{t}$. However, solving $\alpha_{t}$ such that $\alpha_{t}^{\prime}=\alpha_{t}-1$ yields $\alpha_{t}=ce^{t}+1$ and there is no $c$ that satisfies both $\alpha_{0}=1$ and $\alpha_{1}=0$. This shows that the MaskGIT loss (12) may not be a upper bound on the negative log-likelihood. For the linear schedule $\alpha_{t}=1-t$, the backward transition rate matrix conditioned on $x_{0}$ is:
This is the same as the conditional backward transition rate used in Campbell et al. [37, Eq. (22)]—note that their time $t$ is reversed, and the rate matrix was therefore in the form $R^{x_{0}}(t)=\frac{1}{1-t}e_{m}(x_{0}-e_{m})^{\top}$. Please refer to Appendix F for more related work.

### v2

@@ Abstract
Masked (or absorbing) diffusion is actively explored as an alternative to autoregressive models for generative modeling of discrete data. However, existing work in this area has been hindered by unnecessarily complex model formulations and unclear relationships between different perspectives, leading to suboptimal parameterization, training objectives, and ad hoc adjustments to counteract these issues. In this work, we aim to provide a simple and general framework that unlocks the full potential of masked diffusion models. We show that the continuous-time variational objective of masked diffusion models is a simple weighted integral of cross-entropy losses. Our framework also enables training generalized masked diffusion models with state-dependent masking schedules. When evaluated by perplexity, our models trained on OpenWebText surpass prior diffusion language models at GPT-2 scale and demonstrate superior performance on 4 out of 5 zero-shot language modeling tasks. Furthermore, our models vastly outperform previous discrete diffusion models on pixel-level image modeling, achieving 2.75 (CIFAR-10) and 3.40 (ImageNet 64$\times$64) bits per dimension that are better than autoregressive models of similar sizes.
@@ 1 Introduction
  • Using elementary arguments, we establish several properties for the forward process induced by this model and its corresponding time reversal, improving our understanding of this model class.
@@
  • Finally, based on our simplified masked diffusion formulation, we propose a generalized masked diffusion model that allows state-dependent masking schedules. This generalized masked diffusion model further improves predictive performance measured by test likelihoods.
Concurrent work by Ou et al. [36] and Sahoo et al. [37] derives a similar simplified expression of the ELBO. Ou et al. [36]’s derivation relies on an observation similar to the one we made in Proposition 1.
@@ 2 Masked Diffusion
The forward process is defined as a Markovian sequence of discrete random variables $x_{t}$ indexed by time $t$, where $t$ runs from 0 to 1. Throughout the work, we abuse the notation such that $x_{t}$ can be either an integer or its corresponding one-hot vector, whenever it is clear from the context. We divide $[0,1]$ into $T$ intervals, and let $s(i)=(i-1)/T$, $t(i)=i/T$. Following Austin et al. [14], the state transition between $[s(i),t(i)]$ is determined by a transition matrix of size $(m+1)\times(m+1)$: $Q_{i}=(1-\beta_{i})I+\beta_{i}\mathbf{1}e_{m}^{\top},$ where $\mathbf{1}$ is an all-one vector of size $m+1$, $e_{m}$ represents a one-hot vector where element at index $m$ is 1. Each entry $[Q_{i}]_{jk}$ denotes the probability of transition from the state $j$ to the state $k$:
@@
This means that, with probability $1-\beta_{i}$, $x_{t(i)}=x_{s(i)}$, otherwise it jumps to the mask state. Given the above transition matrix, the marginal distribution at time $t(i)$ given $x_{0}$ is
@@
We can define a continuous-time forward process by taking a limit of the above discrete-time process. We first specify a continuous function $\beta(t)$ such that $\beta_{i}=\beta(t(i))/T$. We then let $T\to\infty$ in the discrete-time process and compute the limit of $\bar{Q}_{i}$ (proved in Austin et al. 14, Appendix A.6, see also App. A) as
@@
Figure 1: Masking schedules in the literature: (Left) $\alpha_{t}$; (Right) weight of the cross-entropy loss w.r.t. $t$; Equations for these schedules are given in Tab. 4 in Appendix.
@@
From the definition of $\alpha_{t}$, we have that $\alpha_{0}=1$. And similar to the discrete-time formulation, we would like $\alpha_{1}$ be zero or very close to zero. We provide a summary of masking schedules from literature that satisfy these properties in Fig. 1. The linear schedule was proposed in Sohl-Dickstein et al. [1] for binary variables and then re-derived by Austin et al. [14] from mutual information for discrete-time models. The geometric schedule $\alpha_{t}$ is plotted for $\bar{\beta}_{\text{min}}=10^{-5}$ and $\bar{\beta}_{\text{max}}=20$. It was first used for continuous diffusions [3] and then for discrete by Lou et al. [32]. The cosine schedule was originally proposed in MaskGIT [39], an iterative unmasking generative model inspired by diffusion. This schedule has the property of slowing down the unmasking process at the beginning of the reverse generation. Aligning with their observation, we find that this results in a lower chance of conflicting tokens being unmasked simultaneously at the start of generation, thereby enhancing the overall generation quality.
@@
The analytic property of our forward process allows to compute many quantities of interest in closed form. One such quantity frequently used in diffusion models is the time reversal of the forward process given $x_{0}$: $q(x_{s}|x_{t},x_{0})$ for $s\leq t$. We derive it in App. C as
@@ 3 Model and Objective
For a discrete-time masked diffusion process, we define our generative model by approximately reversing the forward transitions using a reverse model $p_{\theta}(x_{s}|x_{t})$. One way to define this model is
@@
This is known as mean-parameterization since it leverages a prediction model for the mean of $x_{0}$. A matrix-form depiction of $p_{\theta}(x_{s}|x_{t})$ is shown in Fig. 6 (right). In fact, we can select a time-invariant parametrization $\mu_{\theta}(x_{t},t)=\mu_{\theta}(x_{t})$ as [36] showed that $p(x_{0}|x_{t})$ given $x_{t}=x$ is identical for any $t$.
Besides $p_{\theta}(x_{s}|x_{t})$, we also need to specify $p(x_{0}|x_{t(1)})$ and the prior distribution $p(x_{t(T)})=p(x_{1})$. Following the practice in continuous diffusion models [33], we choose $p(x_{0}|x_{t(1)})\propto q(x_{t(1)}|x_{0})$. And since $q(x_{1}|x_{0})\approx\delta_{x_{1},m}$ for any $x_{0}$ as $\alpha_{1}\approx 0$, we set $p(x_{1})\approx\delta_{x_{1},m}$, see App. E.
@@
where $\mathcal{L}_{T}=\sum_{i=2}^{T}\mathbb{E}_{q(x_{t(i)}|x_{0})}[\mathrm{KL}({q(x_{s(i)}|x_{t(i)},x_{0})}\|{p_{\theta}(x_{s(i)}|x_{t(i)})})]$. For the above choices of the prior distribution, the term $\mathrm{KL}({q(x_{1}|x_{0})}\|{p(x_{1})})$ becomes zero. Under the reverse model (2), the KL divergence terms in $\mathcal{L}_{T}$ becomes (proof in App. D)
@@
which is a simple cross-entropy loss between the predicted logits and the clean data. In App. D, we show that $\mathcal{L}_{T}$ is a Riemann sum and is lower bounded by the corresponding continuous integral:
@@
For continuous state-space diffusions, the ELBO depends on the signal-to-noise ratio (SNR) at its endpoints but is otherwise invariant to the noise schedule [33]. We establish here a similar result for discrete diffusions. Consider choosing $\alpha_{t}=\sigma(\lambda_{t})$, where $\sigma$ represents the sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$. In this context, the log-SNR is defined by $\lambda_{t}=\log\frac{\alpha_{t}}{1-\alpha_{t}}=\textup{log-SNR}(t)$. By making a change of variables in (4) to make everything a function of the log-SNR, we obtain
$\mathcal{L}_{\infty}=\int_{\lambda_{t(1)}}^{\lambda_{1}}\sigma(\lambda)\mathbb{E}_{\tilde{q}(x_{\lambda}|x_{0})}\left[\delta_{x_{\lambda},m}\cdot x_{0}^{\top}\log\tilde{\mu}_{\theta}(x_{\lambda},\lambda)\right]\mathrm{d}\lambda.$
@@
In the previous sections, $x_{t}$ was assumed to be a single discrete token. To extend the method to multidimensional data, let $x_{t}$ be now a sequence $(x_{t}^{(1)},x_{t}^{(2)},\dots,x_{t}^{(N)})$, where each element $x_{t}^{(n)}$ represents a discrete token. We select a forward process which factorizes across all $N$ tokens: $q(x_{t}|x_{s})=\prod_{n=1}^{N}q(x_{t}^{(n)}|x_{s}^{(n)})$. As a result, the forward marginals $q(x_{t}|x_{0})$ and reversal $q(x_{s}|x_{t},x_{0})$ also factorize. In this case, we define the reverse model as $p_{\theta}(x_{s}|x_{t})\triangleq\prod_{n=1}^{N}q(x_{s}^{(n)}|x_{t}^{(n)},\mu_{\theta}^{(n)}(x_{t},t))$, where $\mu_{\theta}(x_{t},t)$ is a neural network that takes the full $N$ tokens as input and outputs $N$ probability vectors. [^2] We intentionally choose the reverse model to factorize across dimensions because the true reverse transition $q(x_{s}|x_{t})$ factorizes in the continuous-time limit (as $s$ approaches $t$). The $n$-th output $\mu_{\theta}^{(n)}(x_{t},t)$ is a prediction model for $\mathbb{E}[x_{0}^{(n)}|x_{t}]$, the mean value of the $n$-th token. Repeating above derivations gives
@@
A single step of our training algorithm is described in Alg. 1 in Appendix.
4 Sampling
We use ancestral sampling from our discrete-time reverse process for generation. We have found this yields slightly higher sample quality compared to other methods such as Euler discretization [29, 32]. For conditional generation tasks such as infilling, we find that the simple approach works best — we keep the conditioning tokens unmasked throughout the generation process. A complete description of the sampling algorithm can be found in Alg. 2 in Appendix.
Impact of schedules and discretization.
For comparing different sampling configurations, we primarily use the FID score [40] on image datasets as our evaluation metric. We favor it over text generative perplexity [^3] Perplexity of generated samples scored by a large language model such as GPT-2. used in prior work [32], as the latter can be misleadingly reduced by lowering sample diversity [41]. We initially trained our model using the linear schedule, which achieves the best final ELBO overall; however, we found that sampling did not perform well with a standard uniform discretization grid $t(i)=\frac{i}{T}$. We hypothesize that time discretization can lead to conflicts by generating multiple tokens in a single step. We then switched to the cosine schedule (Tab. 4) that slows down unmasking at the beginning of reverse process. This drastically improves the FID on ImageNet 64$\times$64 from 70 to 17 for $T=256$ steps (Fig. 2, left). Building on this observation, we suggest using a “cosine” discretization grid for sampling in models trained with a linear schedule:
$t(i)=\cos\Big{(}\frac{\pi}{2}\big{(}1-\frac{i}{T}\big{)}\Big{)}.$ (6)
This induces the same discretization in $\alpha_{t}$ as the cosine schedule with a uniform grid, leading to comparable sample quality, as shown in Fig. 2 (left). In Fig. 2 (right), we plot the number of tokens unmasked per step for linear and cosine schedules with a uniform grid. We believe the cosine schedule performs better because it leverages information redundancy: with more tokens revealed, the remaining tokens become more predictable, reducing conflicts when unmasking them in a single step.
Figure 2: Left: FID evaluation for 50K samples randomly generated from our model on pixel-level modeling of ImageNet 64$\times$64. Right: Number of tokens revealed in each step of generation.
Although these findings were originally developed on images, we find them translate well to text (see Fig. 9). we expect other techniques such as top-$p$ sampling [41], classifier-free guidance [42, 43], and predictor-correctors [29, 44] to further improve sample quality of our models. While we reserve these for future work, we note that the JAX [45] implementation of categorical sampling implicitly truncates small probabilities, creating a similar effect to top-$p$ sampling. See App. G for details.
@@ 5 Relation to Existing Work
To show the connection with the CTMC view presented in Austin et al. [14], Campbell et al. [29], we can write out the forward and reverse masked diffusion using CTMC machinery. To see this, for a short time $\Delta t$, given $x_{0}$, the Taylor expansions of our forward and reverse transition matrices at $t$ are
@@
$\bar{R}^{x_{0}}(t,t-\Delta t)$ $=I+R^{x_{0}}(t)\Delta t+o(\Delta t)\text{\quad for\quad}R^{x_{0}}(t)\triangleq-\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top},$ (8)
where $Q(t)$ and $R^{x_{0}}(t)$ are known as the transition rate matrices. Austin et al. [14] derived the same $Q(t)$ in App. A.6 of their paper. However, they did not explore the reverse process or a continuous-time objective. Campbell et al. [29] established an alternative expression for the ELBO using the rate matrices. In Sec. H.1, we show how to recover their expression by separating out a constant from our ELBO expression (4) and applying a discrete “integration-by-part”. A key limitation of their expression is that it needs $N$ evaluations of the prediction model $\mu_{\theta}(\cdot,t)$ to compute an inner summation. To circumvent this computational burden, they used a doubly stochastic estimate. However, this leads to significantly higher variance compared to the analytic cross-entropy (4) which only requires one pass of $\mu_{\theta}(\cdot,t)$. Please refer to Sec. H.2 for more details.
@@
While so far we used a prediction model $\mu_{\theta}(x_{t},t)$ for the mean of clean data given $x_{t}$ (i.e., mean parameterization), one can choose other ways of parameterizing the reverse model. Benton et al. [35], Lou et al. [32] proposed to parameterize the discrete “score” $s(x_{t},t)_{j}\triangleq\frac{q_{t}(j)}{q_{t}(x_{t})}$ and introduced a score-based loss for discrete diffusions. In Sec. H.3, we provide an alternative derivation of their loss which is simpler. We show the link between score and mean parameterizations through the following proposition.
Proposition 1 (Score Parameterization vs. Mean Parameterization).
Let $q_{t}$ be the marginal distribution of the masked diffusion defined in Sec. 2 at time $t$. The discrete score $s(x_{t},t)_{j}=\frac{q_{t}(j)}{q_{t}(x_{t})}$ for a mask state $x_{t}=m$ and $j\neq m$ can be expressed as
@@
Proposition 1 (proved in Sec. H.3) implies that a reasonable score model for a mask state is
$s_{\theta}(m,t)_{j}=\frac{\alpha_{t}}{1-\alpha_{t}}\mu_{\theta}(m,t)_{j}.$ (10)
Indeed, substituting (10) into the score-based loss of Lou et al. [32], Benton et al. [35] recovers our objective (4). In Lou et al. [32], the score is parameterized as a neural network without enforcing the constraint in (9). This means the learned reverse model can be incompatible with the forward process. We find that our parameterization, which enforces the constraint, leads to more stable training and better results.
Any-order autoregressive models.
The continuous-time reverse process of our masked diffusion model can be viewed as an any-order autoregressive model (AO-ARM) [46]. To see this, we reorder the tokens according to the timing of their unmasking events in the reverse process. For all tokens, the cumulative distribution functions (CDFs) of unmasking times $\{\tau_{n}\}_{n=1}^{N}$ are identical and satisfy $P(\tau_{n}\leq t)=P(x_{t}^{(n)}=m)=1-\alpha_{t}$. As a result, the ordering is uniformly random across all possible arrangements, and the token prediction during each unmasking event represents a prediction step in AO-ARMs. This connection was initially pointed out in Hoogeboom et al. [47, App. C]. The relation between our simplified ELBO (5) and the AO-ARM objective is independently clarified by Ou et al. [36]. Despite this equivalence, our work demonstrates that the masking schedule $\alpha_{t}$ introduces a new degree of freedom in the design of such models. Variations in $\alpha_{t}$ can lead to different distributions of unmasking times, significantly impacting performance in diffusion-style parallel sampling under time discretization, as shown in Fig. 2.
@@
Due to space constraint, we defer the discussion on other related work, including MaskGIT [39], discrete flow matching [48], SDDM [30], Blackout diffusion [49] and SUNDAE [50], to Sec. H.4.